# Conditional success of adaptive therapy: the role of treatment-pausing thresholds revealed by mathematical modeling SUMMARY

### v1

@@
Conditional success of adaptive therapy: the role of treatment-pausing thresholds revealed by mathematical modeling
@@
Adaptive therapy (AT) improves cancer treatment by controlling the competition between sensitive and resistant cells via treatment holidays. This study highlights the critical role of treatment-pausing thresholds in AT for tumors composed of drug-sensitive and resistant cells. Using a Lotka-Volterra model, the research compares AT with maximum tolerated dose therapy and intermittent therapy, showing that ATβs success largely depends on the threshold at which treatment is paused and resumed, as well as the competition between sensitive and resistant cells. Three scenarios of comparison between AT and others are identified: uniform-decline, conditional-improve, and uniform-improve, illustrating that optimizing the treatment-pausing threshold is crucial for AT effectiveness. Tumor composition, including initial tumor burden and the proportion of resistant cells, influences outcomes. Adjusting threshold values enables AT to suppress resistant subclones, preserving sensitive cells, ultimately improving progression-free survival. These findings underscore the importance of personalized strategies for cancer management and enhancing long-term therapeutic outcomes.
@@
Adaptive therapy, treatment-holiday threshold, cancer dynamics, mathematical modeling, treatment outcomes
@@ INTRODUCTION
Cancer remains one of the leading causes of death worldwide 1, with its biological complexity and heterogeneity posing significant challenges to conventional therapies such as chemotherapy, radiotherapy, and even immunotherapy 2, 3, 4. Therapeutic resistance, a central obstacle in cancer treatment, arises from the therapeutic selection pressure of pre-existing or induced resistant tumor cells 5, 6, 7, 8, 9. This process drives the evolution of tumor tissue, ultimately leading to the dominance of primary or induced refractory subpopulations, and subsequent relapse and treatment failure 10, 11, 12. The inherent heterogeneity and plasticity of tumor cells often render complete eradication exceedingly difficult, highlighting the necessity of strategies that address the entire tumor population and its evolutionary dynamics rather than selectively targeting certain subpopulations4, 13, 14.
Recently, ecological theory has provided an innovative framework for understanding tumor progression, conceptualizing tumors as ecosystems where drug-sensitive and resistant cells compete for resources 15, 16, 17, 18, 19, 20, 21, 22, 23. Within this ecological framework, cancer cells under drug treatment activate regenerative or protective mechanisms to survive near-lethal stress, enabling them to enter a stem cell-like state and transmit resistance to their progeny 24, 25. This adaptation imposes costs, such as reduced proliferation rates, decreased environmental carrying capacity, and intensified interclonal competition 26, 27.
@@
Adaptive therapy (AT), a dynamic treatment strategy, addresses this challenge by adjusting therapeutic intensity and scheduling in response to real-time tumor burden 16, 34, 29. Unlike fixed-dose regimens, AT employs a feedback-controlled approach: treatment is paused when tumor burden declines to certain lower threshold and resumed upon regrowth to an upper threshold. By retaining a reservoir of sensitive cells, this intermittent dosing modality suppresses resistant subclone expansion and prolongs progression-free survival, as evidenced by clinical successes in prostate cancer and other malignancies.
@@
Despite significant advancements 31, 26, 35, 36, 37, further research is needed to refine the optimization of treatment pause thresholds, a core parameter governing AT efficacy, as most existing studies focus on the efficiency of adaptive therapy under predefined thresholds. Several studies have explored the effects of varying thresholds on AT outcomes across different scenarios 38, 37, 39, 36, 40, 41, providing valuable insights that underscore the importance of threshold adjustments and establishing a solid foundation for further investigation. However, a deeper understanding of the dynamic mechanisms driving context-dependent outcomes under heterogeneous physiological conditions remains essential. A systematic analysis of how different thresholds will give rise to distinct effects and the underlying dynamical mechanisms, which in turn will enhance conventional empirical threshold selection to achieve an optimal balance between tumor suppression and resistance mitigation. Moreover, it can further facilitate the development of personalized strategies tailored to different conditions. Consequently, quantitative modeling of threshold-driven tumor evolutionary dynamics is essential for advancing precision in adaptive therapy design.
@@
Therefore, this study leverages a mathematical model of tumor evolutionary dynamics to analyze and explain the effects of threshold on the outcomes of AT under various conditions of competitive interactions between sensitive and resistant subpopulations, elucidating the biophysical principles underlying threshold-mediated resistance control. The findings are expected to provide theoretical validation for personalized adaptive therapy protocols, facilitating a paradigm shift toward dynamically optimized cancer management.
@@
In summary, this study, as a state-of-the-art work, systematically investigates the impact of treatment pause thresholds on adaptive therapy outcomes using a LotkaβVolterra model that stimulates the competitive dynamics between drug-sensitive and drug-resistant tumor cell populations. By quantifying the relationship between treatment-pausing threshold and employing time to progression (TTP) as an indicator of PFS, the analysis demonstrates that AT efficacy critically depends on the competition between two subpopulations of cancer cells: while weak competition limits the benefits of AT, strong competition can even allow for the indefinite delay of disease progression. Furthermore, this work also reveals that AT is not uniformly superior to MTD, as its efficiency depends on tumor composition (initial conditions) and patientsβ personal conditions (parameters). By comparing the PFS by AT and MTD, it identifies three distinct scenarios, 1) that the which are that AT uniformly improves the PFS than MTD regardless of thresholds, 2) that AT has only conditionally improved the PFS and 3) that AT has uniformly worsened the PFS. These findings highlight the necessity of precise threshold tuning to optimize treatment outcomes. Collectively, these findings provide a theoretical foundation for refining empirical threshold selection and advancing personalized adaptive therapy protocols, thereby paving the way for dynamically optimized cancer management.
@@ RESULTS
To investigate the effects of different treatment strategies, the tumor dynamics under therapeutic conditions are modelled by the widely used Lotka-Volterra (LV)-type equations 42, 18, 36, 43, 44, 45, 46 because the cancer cells behave more similar to unicellular organism rather than normal human cells 17, 47, 48, 49. The cancer cells are further assumed to be classified into two types: drug-sensitive cells (), which can be killed by drugs, and drug-resistant cells (), which cannot be eliminated by drugs. These two cell types compete within the cancer microenvironment (Figure 1A), and therefore the LV model is employed as the following ordinary differential equations (ODEs) (1)-(3).
where denotes the dose curve of therapeutic agents and and represent the burdens of the two types of cells respectively. Equation (1) describes the growth and death of sensitive cells, which means that in the absence of therapy, their proliferation is regulated by intrinsic rate , logistic competition and the death rate is . The drug-dependent suppression of sensitive cells is represented by the factor where is the killing strength of therapeutic agents on sensitive cells. Although this factor is multiplied with the proliferation terms, it is the integrated effective term taking both suppression of proliferation and increasing death rate for simplicity. Equation (2) governs resistant cells, which grow logistically at rate and die at constant rate , unaffected by the drug. is the carrying capacity of the cancer microenvironment, indicating the maximum number of cells that can be supported.
@@
As continuous high-intensity dosage leads to the development of therapeutic resistance through Darwinian selection pressures 29, as well as that high dosage of therapeutic agents causes severe toxicity, the therapy holiday, intermittent therapy regimens incorporating treatment holidays were proposed, which strategically suspends drug administration during predefined recovery periods and emerged as a clinically utilized strategy for balancing therapeutic efficacy with toxicity management 43. The IT strategies can be characterized as deterministic treatment cycles as (5).
@@
As described before, AT dynamically adjusts cancer treatment based on tumor response to prolong efficacy and manage resistance. Two primary approaches within adaptive therapy are dose-skipping and dose-modulating regimens. In dose-skipping adaptive therapy, treatment is administered at the maximum tolerated dose until the tumor burden shrinks to a predetermined threshold (), for example, 50% of the baseline burden; therapy is then paused, allowing the tumor to regrow to a specific threshold (), usually, 100% of the baseline, before resuming treatment. In dose-modulating adaptive therapy, treatment doses are adjusted at regular intervals based on tumor response: increasing the dose if the tumor grows and decreasing it if the tumor shrinks. This strategy seeks to tailor therapy dynamically to tumor behavior, potentially reducing toxicity and managing resistance more effectively. Both approaches exploit the competitive dynamics between treatment-sensitive and resistant cancer cells to enhance long-term treatment outcomes and have already succeeded in trials or experiments 35, 29, 55, 56, 31. Based on the above description, the two AT regimens can be formulated as following equations (6) and (7) respectively.
@@
where is the initial tumor burden and is the derivative of representing whether the tumor burden is growing or shrinking. The threholds satisfies and . Notably, the dosing strategies (6) and (7) represent state-dependent switching mechanisms rather than closed-form mathematical definitions. This implementation induces history-denpendent therapeutic decision and consequently, the governing equations (1)-(3) constitute delay differential equations with history-dependent terms .
As a state-of-the-art investigation, this study provides comprehensive parametric analysis rather than case-specific parameter estimation. The parameter ranges derive from established experimental and clinical researches 35, 36, 57, 58, 59. The proliferation rates of sensitive and resistant cells, i.e. and , are derived from in vitro experimental data 35, while the drug-induced killing strength is sourced from the estimation based on several expriments of cell lines57. Given the critical clinical risk associated with large tumor burdens, the initial tumor burden in this study is taken up to 75% of the environmental carrying capacity 58. The initial proportion of resistant cells follows exsited evolutionary dynamical models 59. All the baseline parameter values are summarized in Table 1.
@@
Proliferation rate of sensitive cells | 35 | ||
Proliferation rate of resistant cells | 35 | ||
Death rate of sensitive cells | |||
Death rate of resistant cells | |||
The killing strength of therapeutic agents | 57 | ||
@@
Initial tumor burden | 58 | ||
Initial proportion of resistant cells | 59 |
The treatment-pausing threshold is crucial for the efficiency of adaptive therapy
@@
To investigate the effect of treatment-pausing thresholds on the outcomes of adaptive therapy, a simplified assumption of neutral competition () is considered. The outcomes are primarily characterized by the time to disease progression (TTP), which is typically defined as the time at which the tumor burden grows to a certain percentage greater than its initial burden35, 36, 57, 58, 59. For simplicity, this percentage is set to % in this study.
@@
As illustrated in the example of tumor evolution under adaptive therapy (FigureΒ 2A), during the initial phase of adaptive therapy, drugs are administered to suppress tumor growth primarily through the depletion of sensitive cells. During this stage, the proportion of sensitive cells is high, which aligns with real-world cases. Following the principles of adaptive therapy, treatment is paused when the tumor burden reaches a predefined treatment-pausing threshold , conventionally set at % of the initial tumor burden, i.e., . The pause allows sensitive cells to repopulate, leading to tumor regrowth. When the tumor burden returns to the initial level , treatment resumes. The treatment-pausing threshold thus defines a cycle, from the start of treatment to its suspension. These cycles continue until resistant cells dominate the tumor burden, ultimately leading to disease progression.
@@
To further elucidate this phenomenon, phase-plane analysis reveals distinct dynamical regimes with and without treatment (FigureΒ 2B1-B2). In untreated systems, tumor populations evolve toward coexistence equilibria, as indicated by the nullcline, which mathematically corresponds to the diagonal in FigureΒ 2B1. In the presence of treatment (FigureΒ 2B2), the long-term steady state shifts to the upper-left corner, where resistant cells dominate the tumor, and sensitive cells eventually disappear. From a dynamical perspective, treatment exerts selective pressure, pushing the tumor toward a resistant cell-dominant state. Therefore, the significance of treatment pauses lies in allowing the tumor to remain longer in a state dominated by sensitive cells, thereby delaying the transition to resistance. In this regard, increasing the treatment-pausing threshold can slow the transition of tumor composition to a resistant state, which implies it is possible to improve adaptive therapy outcomes by modulating the threshold and preventing continuous treatment from driving the tumor toward resistance via Darwinian selection 16.
@@
Therefore, the effects of varying the treatment-pausing threshold are examined under different initial tumor states, characterized by the tumor burden () and the fraction of resistant cells () (ref. TableΒ 1). As shown in FiguresΒ 2C1-C3, altering the treatment-pausing threshold significantly impacted TTP. Notably, when the threshold is too low, meaning the tumor never reached the treatment pause threshold even during the initial cycle, adaptive therapy lacks treatment holidays and effectively becomes equivalent to MTD therapy. Consequently, as is round , TTP of AT is de facto that of MTD therapy, as shown in the leftmost segments of FiguresΒ 2C1-C3.
@@
Importantly, for different initial tumor states, three distinct scenarios emerged when comparing adaptive therapy with MTD. The first scenario, uniform-decline (FigureΒ 2C1), occurs when adaptive therapy consistently performs worse than MTD, regardless of the threshold. This scenario typically arises in tumors with a low initial tumor burden. The second scenario, conditionally-improve (FigureΒ 2C2), occurs when adaptive therapy outperforms MTD only at specific threshold values, suggesting that careful threshold selection is required. The third scenario, uniform-improve (FigureΒ 2C3), occurs when adaptive therapy is always superior to MTD.
@@
These findings highlight the complex relationship between the treatment-pausing threshold and TTP. Moreover, as shown in FiguresΒ 2C2-C3, the effect of the threshold is neither monotonic nor continuous, exhibiting discontinuities as the number of treatment cycles changes discretely (Figure S1). Specifically, in tumors with a low initial burden (FiguresΒ 2C1 and C2), increasing the number of treatment pauses reduces drug exposure, which in turn weakens tumor suppression because under these conditions, the initial tumor burden is low, and therefore, environmental resources remain abundant which makes the changes in the burden of sensitive cells have little impact on the growth rate of resistant cells. In contrast, for tumors with a higher initial burden (FigureΒ 2C3), where environmental resources are constrained, fluctuations in sensitive cell numbers significantly affect the growth of resistant cells. In such cases, increasing the number of treatment pauses enables sensitive cells to grow and exert stronger suppression on resistant cells, leading to a marked increase in TTP. This aligns with clinical observations in prostate cancer, where intermittent dosing preserved therapeutic sensitivity 35.
@@
To further investigate the impact of initial tumor states on these three scenarios, we examined how different initial tumor burdens and initial resistant cell fractions influence adaptive therapy outcomes (FigureΒ 2D). When the initial tumor burden exceeds , adaptive therapy uniformly improves TTP, regardless of the initial proportion of resistant cells. However, when the initial tumor burden is below , adaptive therapy can only conditionally improve TTP, implying that careful selection of the treatment-pausing threshold is crucial. In the worst scenario, when the initial tumor burden is very low and the proportion of resistant cells is relatively high (FigureΒ 2D, top-left), adaptive therapy fails to improve TTP at any threshold, indicating primary resistance. These findings extend prior theoretical work 42 by quantifying phase-transition-like phenomena in threshold-mediated outcomes.
@@
Furthermore, clinical implementation faces inherent uncertainty in tumor state assessment and threshold selection, implying that both the treatment-pausing threshold and the initial tumor burden and resistance fraction exhibit random variability. To address this, Monte Carlo simulations were performed to evaluate the probabilities of the three outcome scenarios under parameter variability (FigureΒ 2E). Inspired by the Norton-Simon Hypothesis 60, 61, which posits that sensitive cells typically have a greater fitness advantage than resistant cells (i.e., sensitive cells have a higher intrinsic proliferation rate than resistant cells), we further explored the effect of the relative growth rate of sensitive cells () on the probabilities of the three scenarios. By randomly varying the treatment threshold and, while keeping fixed, randomly altering and , we found that the probability of the uniform-improve scenario increases as the relative fitness of sensitive cells increases. When is small, adaptive therapy is less likely to improve TTP relative to MTD. Conversely, when is high, indicating that sensitive cells have a significant growth advantage over resistant cells, adaptive therapy is more likely to enhance TTP. These results confirm that the treatment-pausing threshold has a profound impact on adaptive therapy outcomes and patient-specific tumor characteristics, including initial tumor burden and the relative fitness of sensitive and resistant cells, significantly influence treatment outcomes. These findings align with existing research 26 and demonstrate that dependence provides quantitative criteria for patient stratification, providing insights for addressing a key challenge in empirical threshold selection 31.
Adaptive therapy is Comparison of the optimal outcomes of four therapeutic strategies
@@
For intermittent therapy, the treatment cycle period () is examined to range from to time units, with drug administration lasting from to time units () per cycle. Since this study does not focus on identifying these optimal parameters of IT, computational details are omitted. The result shows that the suboptimal period is with , simulating a regimen in which the patient undergoes both treatment and rest periods within a feasible schedule (TableΒ 1 and equationsΒ (5)). For dose-skipping adaptive therapy, the treatment-pausing threshold is set to , while for dose-modulating adaptive therapy, and are selected as the suboptimal treatment parameters (FigureΒ 3A1-A4).

### v2

# Conditional Success of Adaptive Therapy: The Role of Treatment-Holiday Thresholds Revealed by Mathematical Modeling SUMMARY

@@
Conditional Success of Adaptive Therapy: The Role of Treatment-Holiday Thresholds Revealed by Mathematical Modeling
@@
Adaptive therapy (AT) improves cancer treatment by controlling the competition between sensitive and resistant cells through treatment holidays. This study highlights the critical role of treatment-holiday thresholds in AT for tumors composed of drug-sensitive and resistant cells. Using a Lotka-Volterra model, the research compares AT with maximum tolerated dose therapy and intermittent therapy, showing that ATβs success largely depends on the threshold at which treatment is paused and resumed, as well as on the competition between sensitive and resistant cells. Three scenarios of comparison between AT and other therapies are identified: uniform-decline, conditional-improve, and uniform-improve, illustrating that optimizing the treatment-holiday threshold is crucial for AT effectiveness. Tumor composition, including initial tumor burden and the proportion of resistant cells, influences outcomes. Adjusting threshold values enables AT to suppress resistant subclones, preserving sensitive cells, ultimately improving progression-free survival. These findings emphasize the importance of personalized treatment strategies potentially enhancing long-term therapeutic outcomes.
@@
adaptive therapy, treatment-holiday threshold, cancer dynamics, mathematical modeling, cell competition
@@ INTRODUCTION
Cancer remains one of the leading causes of death worldwide 1, with its biological complexity and heterogeneity posing significant challenges to conventional therapies such as chemotherapy, radiotherapy, and even immunotherapy 2, 3, 4. Therapeutic resistance, a central obstacle in cancer treatment, arises from the therapeutic selection pressure of pre-existing or induced resistant tumor cells 5, 6, 7, 8, 9. This process drives the evolution of tumor tissue, ultimately leading to the dominance of primary or induced refractory subpopulations, and subsequent relapse and treatment failure 10, 11, 12. The inherent heterogeneity and plasticity of tumor cells often render complete eradication exceedingly difficult, highlighting the necessity of strategies that address the entire tumor population and its evolutionary dynamics, rather than targeting specific subpopulations 4, 13, 14.
Recently, ecological theory has provided an innovative framework for understanding tumor progression, conceptualizing tumors as ecosystems where drug-sensitive and drug-resistant cells compete for resources 15, 16, 17, 18, 19, 20, 21, 22, 23. Within this ecological framework, cancer cells under drug treatment activate regenerative or protective mechanisms to survive near-lethal stress, enabling them to enter a stem cell-like state and transmit resistance to their progeny 24, 25. This adaptation imposes costs, such as reduced proliferation rates, decreased environmental carrying capacity, and intensified interclonal competition 26, 27.
@@
Adaptive therapy (AT) is a dynamic treatment strategy that employs a feedback-controlled approach, adjusting therapeutic intensity and scheduling in response to real-time tumor burden 16, 34, 29: treatment is paused when tumor burden declines to a certain lower threshold and resumed upon regrowth to an upper threshold. FT can be classified into dose-skipping (AT-S) and dose-modulating (AT-M) regimens, depending on whether more than two dose rates are used during treatment phases. By retaining a reservoir of sensitive cells, this intermittent dosing modality suppresses resistant subclone expansion and prolongs progression-free survival, as evidenced by clinical successes in prostate cancer and other malignancies.
@@
Despite significant advancements 31, 26, 35, 36, 37, further research is needed to refine the optimization of treatment-holiday thresholds, a core parameter governing AT efficacy, as most existing studies focus on the efficiency of adaptive therapy under predefined thresholds. Several studies have explored the effects of varying thresholds on AT outcomes across different scenarios 38, 37, 39, 36, 40, 41, providing valuable insights that underscore the importance of threshold adjustments and establishing a solid foundation for further investigation. However, a deeper understanding of the dynamic mechanisms driving context-dependent outcomes under heterogeneous physiological conditions remains essential. It requires a systematic analysis of the underlying dynamical mechanisms by which different thresholds give rise to distinct effects, enhancing conventional empirical threshold selection to achieve an optimal balance between tumor suppression and resistance mitigation. Moreover, it can further facilitate the development of personalized strategies tailored to different conditions. Consequently, quantitative modeling of threshold-driven tumor evolutionary dynamics is essential for advancing precision in adaptive therapy design.
@@
Therefore, this study leverages a mathematical model of tumor evolutionary dynamics to analyze and explain the effects of the thresholds on the outcomes of AT under various conditions of competitive interactions between sensitive and resistant subpopulations, elucidating the biophysical principles underlying threshold-mediated resistance control. The findings are expected to provide theoretical validation for personalized adaptive therapy protocols, facilitating a paradigm shift toward dynamically optimized cancer management.
@@
In summary, this study, as a state-of-the-art work, systematically investigates the impact of treatment-holiday thresholds on adaptive therapy outcomes using a LotkaβVolterra model that simulates the competitive dynamics between drug-sensitive and drug-resistant tumor cell populations. By quantifying the relationship between treatment-holiday threshold and employing time-to-progression (TTP) as an indicator of progression-free survival, the analysis demonstrates that AT efficacy critically depends on the competition between two subpopulations of cancer cells. Strong competition can allow for the indefinite delay of disease progression, while weak competition requires the benefits of AT. Furthermore, this work also reveals that AT is not uniformly superior to MTD, as its efficiency depends on tumor composition (initial conditions) and patientsβ personal conditions (parameters). By comparing the TTP between AT and MTD, three distinct scenarios are identified: 1) uniform-decline, where AT consistently underperforms MTD regardless of the threshold; 2) conditional-improve, where the effectiveness of AT depends on the specific threshold selected; and 3) uniform-improve, where AT consistently outperforms MTD. These findings highlight the necessity of precise threshold tuning to optimize treatment outcomes. Collectively, these findings provide a theoretical foundation for refining empirical threshold selection and advancing personalized adaptive therapy protocols, thereby paving the way for dynamically optimized cancer management.
@@ RESULTS
To investigate the effects of different treatment strategies, the tumor dynamics under therapeutic conditions are modeled by the widely used Lotka-Volterra (LV)-type equations 42, 18, 36, 43, 44, 45, 46 because the cancer cells behave more similarly to unicellular organisms than to normal human cells 17, 47, 48, 49. The cancer cells are further assumed to be classified into two types: drug-sensitive cells (), which can be killed by drugs, and drug-resistant cells (), which cannot be killed by drugs. These two cell types compete within the cancer microenvironment (Figure 1A), and therefore the LV model is described as the following differential equations (1)-(3).
where denotes the dose curve of therapeutic agents and and represent the burdens of the two types of cells respectively. Equation (1) describes the growth and death of sensitive cells, which means that in the absence of therapy, their proliferation is regulated by intrinsic rate , logistic competition and the death rate is . The drug-dependent suppression of sensitive cells is represented by the factor where is the killing strength of therapeutic agents on sensitive cells. Although this factor is multiplied with the proliferation terms, it is the integrated effective term taking both suppression of proliferation and increasing death rate for simplicity. Equation (2) governs resistant cells, which obey logistical growth at rate and die at constant rate , unaffected by the drug. is the carrying capacity of the cancer microenvironment, indicating the maximum number of cells that can be supported.
@@
As continuous high-intensity dosage leads to the development of therapeutic resistance through Darwinian selection pressures 29, as well as that high dosage of therapeutic agents causes severe toxicity, the intermittent therapy regimens were proposed, which strategically suspend drug administration during predefined recovery periods and emerged as a clinically utilized strategy for balancing therapeutic efficacy with toxicity management 43. The IT strategies can be characterized as deterministic treatment cycles as (5).
@@
As described before, AT dynamically adjusts cancer treatment based on tumor response to prolong efficacy and manage resistance. Two primary approaches within adaptive therapy are dose-skipping and dose-modulating regimens. In dose-skipping adaptive therapy, treatment is administered at the maximum tolerated dose until the tumor burden shrinks to a predetermined threshold (), for example, % of the baseline burden; therapy is then paused, allowing the tumor to regrow to a specific threshold (), usually, % of the baseline, before resuming treatment. In dose-modulating adaptive therapy, treatment doses are adjusted at regular intervals based on tumor response: increasing the dose if the tumor grows and decreasing it if the tumor shrinks. This strategy seeks to tailor therapy dynamically to tumor behavior, potentially reducing toxicity and managing resistance more effectively. Both approaches exploit the competitive dynamics between treatment-sensitive and resistant cancer cells to enhance long-term treatment outcomes and have already succeeded in trials or experiments 35, 29, 55, 56, 31. Based on the above description, the two AT regimens can be formulated as following equations (6) and (7) respectively.
@@
where is the initial tumor burden and is the derivative of representing whether the tumor burden is growing or shrinking. The dose-modulating thresholds satisfy and . Notably, the dosing strategies (6) and (7) represent state-dependent switching mechanisms rather than closed-form mathematical definitions. This implementation induces history-dependent therapeutic decision and consequently, the governing equations (1)-(3) constitute delay differential equations with history-dependent terms .
As a state-of-the-art investigation, this study provides comprehensive parameter analysis rather than case-specific parameter estimation. So, in this work, the parameter values are taken in ranges according to established experimental and clinical studies 35, 36, 57, 58, 59. The proliferation rates of sensitive and resistant cells, i.e. and , are derived from in vitro experimental data 35, while the drug-induced killing strength is sourced from the estimation based on several experiments of cell lines57. Given the critical clinical risk associated with large tumor burdens, the initial tumor burden in this study is set as 75% of the environmental carrying capacity 58. The initial proportion of resistant cells is taken from existing dynamic models 59. All the baseline parameter values are summarized in Table 1.
@@
| Proliferation rate of sensitive cells | Zhang etΒ al. 35 | ||
| Proliferation rate of resistant cells | Zhang etΒ al. 35 | ||
| Death rate of sensitive cells | Zhang etΒ al. 35 | ||
| Death rate of resistant cells | Zhang etΒ al. 35 | ||
| The killing strength of therapeutic agents | West etΒ al. 57 | ||
@@
| Initial tumor burden | Sunassee etΒ al. 58 | ||
| Initial proportion of resistant cells | Grassberger etΒ al. 59 | ||
| , | The treatment-holiday thresholds for dose-skipping adaptive therapy | ||
| , , | The dose-modulating thresholds for dose-modulating adaptive therapy | ||
The treatment-holiday threshold is crucial for the efficiency of adaptive therapy
@@
To investigate the effect of treatment-holiday thresholds on the outcomes of adaptive therapy, a simplified assumption of neutral competition () is considered. The outcomes are primarily characterized by the time to progression (TTP), which is typically defined as the time at which the tumor burden grows to a certain percentage greater than its initial burden35, 36, 57, 58, 59. For simplicity, this percentage is set to % in this study.
@@
As illustrated in the example of tumor evolution under adaptive therapy (FigureΒ 2A), during the initial phase of adaptive therapy, drugs are administered to suppress tumor growth primarily through the depletion of sensitive cells. During this stage, the proportion of sensitive cells is high, which aligns with real-world cases. Following the principles of adaptive therapy, treatment is paused when the tumor burden reaches a predefined treatment-holiday threshold , conventionally set at % of the initial tumor burden, i.e., . The treatment holidays allow sensitive cells to repopulate, leading to tumor regrowth. When the tumor burden returns to the initial level , treatment resumes. The treatment-holiday threshold thus defines a cycle, from the start of treatment to its suspension. These cycles continue until resistant cells dominate the tumor burden, ultimately leading to disease progression.
@@
To further elucidate this phenomenon, phase-plane analysis reveals distinct dynamical regimes with and without treatment (FigureΒ 2B1-B2). In untreated systems, tumor populations evolve toward coexistence equilibria, as indicated by the nullcline, which mathematically corresponds to the diagonal in FigureΒ 2B1. In the presence of treatment (FigureΒ 2B2), the long-term steady state shifts to the upper-left corner, where resistant cells dominate the tumor, and sensitive cells eventually disappear. From a dynamical perspective, treatment exerts selective pressure, pushing the tumor toward a resistant cell-dominant state. Therefore, the significance of treatment holidays lies in allowing the tumor to remain longer in a state dominated by sensitive cells, thereby delaying the transition to resistance. In this regard, increasing the treatment-holiday threshold can slow the transition of tumor composition to a resistant state, which implies it is possible to improve adaptive therapy outcomes by modulating the threshold and preventing continuous treatment from driving the tumor toward resistance via Darwinian selection 16.
@@
Therefore, the effects of varying the treatment-holiday threshold are examined under different initial tumor states, characterized by the tumor burden () and the fraction of resistant cells () (ref. TableΒ 1). As shown in FiguresΒ 2C1-C3, altering the treatment-holiday threshold significantly impacted TTP. Notably, when the threshold is too low, meaning the tumor never reached the treatment-holiday threshold even during the initial cycle, adaptive therapy lacks treatment holidays and effectively becomes equivalent to MTD therapy. Consequently, as is around , TTP of AT is de facto that of MTD therapy, as shown in the leftmost segments of FiguresΒ 2C1-C3.
@@
Importantly, for different initial tumor states, three distinct scenarios emerged when comparing adaptive therapy with MTD. The first scenario, uniform-decline (FigureΒ 2C1), occurs when adaptive therapy consistently performs worse than MTD, regardless of the threshold. This scenario typically arises in tumors with a low initial tumor burden. The second scenario, conditional-improve (FigureΒ 2C2), occurs when adaptive therapy outperforms MTD only at specific threshold values, suggesting that careful threshold selection is required. The third scenario, uniform-improve (FigureΒ 2C3), occurs when adaptive therapy is always superior to MTD.
@@
The findings highlight the complex relationship between treatment-holiday thresholds and TTP. Moreover, as shown in FiguresΒ 2C2-C3, the effect of the threshold is neither monotonic nor continuous, exhibiting discontinuities as the number of treatment cycles changes discretely (Figure S1). Specifically, in tumors with a low initial burden (FiguresΒ 2C1 and C2), increasing the number of treatment holidays reduces drug exposure, which in turn weakens tumor suppression because under these conditions, the initial tumor burden is low, and therefore, environmental resources remain abundant which makes the changes in the burden of sensitive cells have little impact on the growth rate of resistant cells. In contrast, for tumors with a higher initial burden (FigureΒ 2C3), where environmental resources are constrained, fluctuations in sensitive cell numbers significantly affect the growth of resistant cells. In such cases, increasing the number of treatment holidays enables sensitive cells to grow and exert stronger suppression on resistant cells, leading to a marked increase in TTP. This aligns with clinical observations in prostate cancer, where intermittent dosing preserved therapeutic sensitivity 35.
@@
To further investigate the impact of initial tumor states on these three scenarios, we examined how different initial tumor burdens and initial resistant cell fractions influence adaptive therapy outcomes (FigureΒ 2D). When the initial tumor burden exceeds , adaptive therapy uniformly improves TTP, regardless of the initial proportion of resistant cells. However, when the initial tumor burden is below , adaptive therapy can only conditionally improve TTP, implying that careful selection of the treatment-holiday threshold is crucial. In the worst scenario, when the initial tumor burden is very low and the proportion of resistant cells is relatively high (FigureΒ 2D, top-left), adaptive therapy fails to improve TTP at any threshold, indicating primary resistance. These findings extend prior theoretical work 42 by quantifying phase-transition-like phenomena in threshold-mediated outcomes.
@@
Furthermore, clinical implementation faces inherent uncertainty in tumor state assessment and threshold selection, implying that both the treatment-holiday threshold and the initial tumor burden and resistance fraction exhibit random variability. To address this, Monte Carlo simulations were performed to evaluate the probabilities of the three outcome scenarios under parameter variability (FigureΒ 2E). Inspired by the Norton-Simon Hypothesis 60, 61, which posits that sensitive cells typically have a greater fitness advantage than resistant cells (i.e., sensitive cells have a higher intrinsic proliferation rate than resistant cells), we further explored the effect of the relative growth rate of sensitive cells () on the probabilities of the three scenarios. By randomly varying the treatment threshold and, while keeping fixed, randomly altering and , we found that the probability of the uniform-improve scenario increases as the relative fitness of sensitive cells increases. When is small, adaptive therapy is less likely to improve TTP relative to MTD. Conversely, when is high, indicating that sensitive cells have a significant growth advantage over resistant cells, adaptive therapy is more likely to enhance TTP. These results confirm that the treatment-holiday threshold has a profound impact on adaptive therapy outcomes and patient-specific tumor characteristics, including initial tumor burden and the relative fitness of sensitive and resistant cells, significantly influence treatment outcomes. These findings align with existing research 26 and demonstrate that dependence provides quantitative criteria for patient stratification, providing insights for addressing a key challenge in empirical threshold selection 31.
Context-dependent superiority of adaptive therapy across treatment strategies
@@
For intermittent therapy, the treatment cycle period () is examined to range from to time units, with drug administration lasting from to time units () per cycle. Since this study does not focus on identifying these optimal parameters of IT, computational details are omitted. The result shows that the suboptimal period is with , simulating a regimen in which the patient undergoes both treatment and rest periods within a feasible schedule (TableΒ 1 and equationsΒ (5)). For dose-skipping adaptive therapy, the treatment-holiday threshold is set to , while for dose-modulating adaptive therapy, and are selected as the suboptimal treatment parameters (FigureΒ 3A1-A4).